# Learning Dynamic Bayesian Networks from Data: Foundations, First Principles and Numerical Comparisons

### v1

@@ Abstract
In this paper, we present a guide to the foundations of learning Dynamic Bayesian Networks (DBNs) from data in the form of multiple samples of trajectories for some length of time. We present the formalism for a generic as well as a set of common types of DBNs for particular variable distributions. We present the analytical form of the models, with a comprehensive discussion on the interdependence between structure and weights in a DBN model and their implications for learning. Next, we give a broad overview of learning methods and describe and categorize them based on the most important statistical features, and how they treat the interplay between learning structure and weights. We give the analytical form of the likelihood and Bayesian score functions, emphasizing the distinction from the static case. We discuss functions used in optimization to enforce structural requirements. We briefly discuss more complex extensions and representations. Finally we present a set of comparisons in different settings for various distinct but representative algorithms across the variants.
@@ 1 Introduction
In this paper we give a comprehensive presentation on the training of Dynamic Bayesian Networks (DBNs), including both structure and parameters, from data. DBNs present a naturally interpretable model when it comes to understanding the precise interaction underlying the relationship between the variables. That is, the conditional independence structure defined by the DBN provides information regarding the mechanistic procedure that defines the model. This is also associated with the field of statistics referred to as causal learning.
There is one general survey article on DBNs we found is [59], which provides a helpful comprehensive resource for references for DBN modeling, inference and learning. In this work rather than seeking to provide a comprehensive literature review, instead we focus on narrating the global landscape of the mathematical understanding of the most important considerations as far as learning a model, including both the structure and parameters, from data.
With this understanding we are able to establish an informative taxonomy regarding methods, providing transparency in regards to the function and intention of each method. There are important subtle distinctions as far as modeling assumptions between some different popular methods, and their awareness is critical for best practices of DBN learning. Finally, we perform comprehensive numerical comparisons, highlighting the particular advantages and disadvantages of each method. We highlight that these comparisons are not meant to be exhaustive or authoritative, but more informative and illustrative in regards to the tradeoffs associated with learning DBNs from data.
We make several assumptions for this work. These are not entirely formal for the sake of precise Theorem proving, but rather a general restriction of the data generating process of interest, so as to highlight the most pedagogical features of DBN learning, as far as it is most commonly done in the literature. These assumptions are restrictive as far as faithful statistical modeling of real world phenomena. However, they tremendously simplify the learning process, and thus allow a more comprehensible presentation of what can hope to achieve with standard simple methods. For transparency, we present them below.
@@
    Causal Identifiability we focus on the general case wherein the data regime permits for potential identification of a true causal graph, generally corresponding to the number of data samples (trajectories $\times$ time steps) some exponential factor of mangitude greater than the number of variables (in practice this can mean $5$ variables, $100$ trajectories of each $50$ time steps as a generic example). This permits us to focus on integer programming and other techniques that can obtain statistically significant point estimates for exact structures that recover a ground truth. This assumption is standard in the literature of DBN and SEM learning, as modeling uncertainty in less favorable data regime circumstances presents significant methodological challenges and considerations that require significantly more advanced techniques. However, understanding the nuances of the foundations are essential as far the proper development, implementation, and use of these techniques.
@@
    Fully Observed There are no hidden variables, all quantities of interest are fully observed at every time step. Of course, graphs with hidden (latent) variables and entire structures are instrumental for modeling in many fields. However, the inclusion of latent variables and the required Expectation Maximization modification of the procedures described presents technical complications that add significant additional complexity, and thus would necessity a much greater length while obfuscating the message.
We make a few departures from these assumptions throughout the work, which we explicitly indicate when they occur.
@@ 1.2 Applications of Dynamic Bayesian Networks
The discovery of dynamic Bayesian networks has found many applications, some of which are medicine [12, 20, 75, 8], economics [43, 44] and aviation [46, 65, 27]. The applications related to aviation are typically related to finding casual structure in a sub-problem related to the dispatch of flights and focused on risk mitigation. The medial applications typically focus on either the discovery of fundamental principles related to chemical reactions, which take place in biological organisms or the extraction of information from clinical data. In economics, it is typically of interest to uncover relationships between the stock market and other selected factors that either influence or are influenced by it. In the following, we give some details about chosen applications as well as specific outcomes that the modelling using DBNs has in practice.
@@ 1.2.1 Medical Science
To highlight the importance of DBN discovery in medicine, we detail three separate applications. The first of these focuses on the quantification of disease development [8]. Understanding the progression of diseases is crucial in clinical medicine, as it informs the effectiveness of treatments. Most clinical medicine and pathology textbooks provide detailed descriptions of disease progression and treatment responses. However, there has been limited research quantifying these descriptions in detail. Typically, research examines the temporal aspect by describing treatment outcomes after a certain period. A significant challenge in gaining deeper insights is the relatively small size of clinical datasets, often comprising only a few hundred patients.
In the aforementioned contribution, a heuristic procedure is proposed for exploring and learning non-homogeneous time dynamic Bayesian networks, aiming to balance specificity and simplicity. The approach begins with a fully homogeneous (in time) model parts of which are gradually replaced with sub-models which reflect the expected structure at a given level of time delay. Furthermore, a splitting technique is applied to further improve the predictive behavior of the model, such models are typically termed partitioned DBNs.
@@
The second application is concerned with the mapping of neural pathways [20]. Identifying functional connectivity from simultaneously recorded spike trains is crucial for understanding how the brain processes information and instructs the body to perform complex tasks. The study investigates the applicability of dynamic Bayesian networks (DBNs) to infer the structure of neural circuits from observed spike trains. A probabilistic point process model was employed to assess performance. The results confirm the utility of DBNs in inferring functional connectivity as well as the directions of signal flow in cortical network models. Additionally, the findings demonstrate that DBNs outperform Granger causality when applied to populations with highly non-linear synaptic integration mechanisms.
The third chosen application focuses on the choice of appropriate treatment regimens for Chronic lymphocytic leukemia (CLL) [75]. This cancer is the most common blood cancer in adults, with a varied course and response to treatment among patients. This variability complicates the selection of the most appropriate treatment regimen and the prediction of disease progression. The aforementioned paper aims to develop and validate dynamic Bayesian networks (DBNs) to predict changes in the health status of patients with CLL and predict the progression of the disease over time. Two DBNs, the Health Status Network (HSN) and the Treatment Effect Network (TEN), were developed and implemented. Relationships linking the most important factors influencing health status and treatment effects in CLL patients were identified based on literature data and expert knowledge. The developed networks, particularly TEN, were able to predict the probability of survival in CLL patients, aligning with survival data collected in large medical registries. The networks can tailor predictions by integrating prior knowledge specific to an individual CLL patient. The proposed approach is a suitable foundation for developing artificial intelligence systems that assist in selecting treatments, thereby positively influencing the chances of survival for CLL patients.
@@ 1.2.2 Economics
The relationship between the stock market and national economies deepens as the market matures, highlighting the need to study their dynamic interplay. Economic indicators such as real income and savings rates play crucial roles in influencing stock market capitalization. Macroeconomic fundamentals wield considerable influence over both short and long-term periods. Some researchers argue that finance and economic growth are causally linked, suggesting the stock market’s potential to drive economic development. However, not all macroeconomic factors significantly impact stock prices. Understanding the strength of association among these variables offers insights into how the stock market behaves across varying economic landscapes. Research on the Chinese stock market examines how macroeconomic variables shape stock market indices over time, emphasizing the enduring influence of economic fundamentals amid short-term market fluctuations. The aforementioned interplay may be modeled by DBNs and has been detailed in [44].
@@ 1.2.3 Aviation
The global collection of aircraft and the airspace in which they operate, is a complex system generating a vast amount of data, making it a challenging domain to model mathematically. This system includes critical elements such as aircraft, airports, flight crews, weather events, and routes, each with many subcomponents. For example, aircraft have numerous subsystems and components, each subjected to various stresses and maintenance actions, which influence their time dynamics. Airports have multiple runways and internal logistical processes, which influence the operational capacity. These system components interact in complex ways. For instance, each flight corresponds to an aircraft operated by a flight crew traveling from one airport to another via a route that may need to change due to weather. Multiple flights operate simultaneously, requiring coordination to avoid incidents while maximizing throughput and minimizing delays.
To give an idea about a specific aviation problem that may be tackled we describe the airport operation uncertainty characterisation, which has been developed in [27]. The model outlines aircraft flow through the airport, emphasizing integrated airspace and airside operations. It characterizes various operational milestones based on an aircraft flow’s Business Process Model and Airport Collaborative Decision-Making methodology. Probability distributions for factors influencing aircraft processes need to be estimated, along with their conditional probability relationships. This approach results in a dynamic Bayesian network that manages uncertainties in aircraft operating times at the airport. The nodes of the network describe various aspects of the airport and flight operations. They cover meteorological conditions, arrival airspace variables such as timestamps and congestion metrics, airport infrastructure, operator and flight data, airside operational times and flight regulations, and the causes of delays.
@@ 2 Background - Dynamic Bayesian Networks
We present the general, and then specific forms, of DBN models. Consider that there is an $n$-dimensional stochastic process $X(t)$. The individual random variables $X_{i}(t)$ for all $i\in[n_{x}]$ can be valued as discrete, or as members of some field, such as $\mathbb{R}$. In addition, there can be an $n_{z}$-dimensional random variable $Z$. Let us denote the generic spaces as $\mathcal{X}$ and $\mathcal{Z}$, respectively.
The defining character of DBNs is modeling the dependence of $X_{i}(t)$ on other quantities, i.e., defining the the evolution of the stochastic process $X(t)\to X(t+1)$. Formally, for the probability kernel defining the iterations of the stochastic process, the dependence must be Markovian, that is
@@
In addition, there is the important requirement that no in-time string of dependencies forms a cycle. This presents the necessity in introducing graph theoretical notions to precisely characterize DBNs. Generically, we say a directed graph is a set of vertices $\mathcal{V}=\{v_{1},v_{2},...,v_{n}\}$ and edges =$\mathcal{E}=\{e_{1},e_{2},...,e_{m}\}$, where $e_{j}=(v_{j_{1}},v_{j_{2}})$ denotes the existence of a directed path between the two notes $v_{j_{2}}\to v_{j_{1}}$. We also say in this case that $j_{2}$ is a parent of $j_{1}$, or $j_{2}\in dpa(j_{1})$.
The DBN model incorporates a directed acyclic graph (DAG) $\bar{\mathcal{G}}=\mathcal{G}(\mathcal{V}(X(t-1),X(t)),\mathcal{E}_{d})\cup\mathcal{G}(\mathcal{V}(X(t),Z),\mathcal{E}_{z})\cup\mathcal{G}(\mathcal{V}(X(t),X(t-\tau)),\mathcal{E}_{\tau})$. The first two define connections in the model between the temporal random variables. That is $e=\{V_{1},V_{2}\}\in\mathcal{E}_{d}$ with $V_{i}\in\{\{X_{i}\}\}$ if $p(X(t+1)_{i}\in A)$ is a function of $V_{2}=X_{j}$, that is $j\in dpa_{d}(i)$, and $e=\{V_{1},V_{2}\}\in\mathcal{E}_{s}$ if $p(X_{i}(t+1)\in A)$ is a function of $V_{2}=X_{j}(t+1)$, that is $2\in dpa_{s}(i)$. Finally, we also have a (non-symmetric) matrix encoding the dependencies on the self-history $\tau\in dpa_{\tau}(i)\subset\{1,...,p\}$ and the dependencies on the static variables $dpa_{z}(i)=\left\{Z_{j}:\frac{\partial f_{i}(\cdot)}{\partial Z_{j}}\neq 0\right\}$. These, of course, can be encoded as graphs as well.
@@
See Figure 1 for an illustration. In this case for $i=1$, the Markovian transitions are from itself and from $X_{2}$, and there are no intra time nodes or static nodes directed to it, and so $V_{dpa(1)}(t)=V_{dpa_{d}(1)}=\{X_{1}(t-1),X_{2}(t-1)\}$. For node 2, there are no Markovian transitions and two intra-node dependencies, thus $V_{dpa(2)}(t)=V_{dpa_{s}(2)}(t)=\{X_{1}(t),X_{3}(t)\}$. Finally, for $X_{3}(t)$, there are two Markovian dependencies, and a static covariate dependence. Thus $V_{dpa(3)}(t)=V_{dpa_{d}(3)\cup dpa_{z}(3)}(t)=\{X_{1}(t-1),X_{3}(t-1),Z\}$.

### v2

@@ Abstract
In this paper, we present a guide to the foundations of learning Dynamic Bayesian Networks (DBNs) from data in the form of multiple samples of trajectories for some length of time. We present the formalism for a generic as well as a set of common types of DBNs for particular variable distributions. We present the analytical form of the models, with a comprehensive discussion on the interdependence between structure and weights in a DBN model and their implications for learning. We give the analytical form of the likelihood and Bayesian score functions, emphasizing the distinction from the static case. We discuss functions used in optimization to enforce structural requirements. Next, we give a broad overview of learning methods and describe and categorize them based on the most important statistical features, and how they treat the interplay between learning structure and weights. We briefly discuss more complex extensions and representations. Finally we present a set of comparisons in different settings for various distinct but representative algorithms across the variants.
@@ 1 Introduction
In this paper we give a comprehensive presentation on the training of Dynamic Bayesian Networks (DBNs), including both structure and parameters, from data. DBNs present a naturally interpretable model when it comes to understanding the precise interaction underlying the relationship between random variables. That is, the conditional independence structure defined by the DBN provides information regarding the mechanistic procedure that defines the model. This is also associated with the field of statistics referred to as causal learning.
There is one general survey article on DBNs we found is [shiguihara2021dynamic], which provides a helpful comprehensive resource for references for DBN modeling, inference and learning. In this work rather than seeking to provide a comprehensive literature review, instead we focus on narrating the global landscape of the mathematical understanding of the most important considerations as far as learning a model, including both the structure and parameters, from data.
With this understanding we are able to establish an informative taxonomy regarding methods, providing transparency in regards to the function and intention of each approach. There are important subtle distinctions as far as modeling assumptions between some different popular methods, and their awareness is critical for best practices of DBN learning. Finally, we perform comprehensive numerical comparisons, highlighting the particular advantages and disadvantages of each method. We highlight that these comparisons are not meant to be exhaustive or authoritative, but more informative and illustrative in regards to the tradeoffs associated with learning DBNs from data.
We make several assumptions for this work. These are not entirely formal for the sake of precise Theorem proving, but rather a general restriction of the data generating process of interest, so as to highlight the most pedagogical features of DBN learning, as far as it is most commonly done in the literature. These assumptions are restrictive as far as faithful statistical modeling of real world phenomena. However, they tremendously simplify the learning process, and thus allow a more comprehensible presentation of what can one hope to achieve with standard simple methods. For transparency, we present them below:
@@
    Causal Identifiability we focus on the general case wherein the data regime permits for potential identification of a true causal graph, generally corresponding to the number of data samples (trajectories $\times$ time steps) some exponential factor of mangitude greater than the number of variables (in practice this can mean $5$ variables, $100$ trajectories of each $50$ time steps as a generic example). This permits us to focus on integer programming, constraint based methods, and other techniques that can obtain statistically significant point estimates for exact structures that recover a ground truth. This assumption is standard in the literature of DBN and SEM learning, as modeling uncertainty in less favorable data regime circumstances presents significant methodological challenges and considerations that require significantly more advanced techniques. However, understanding the nuances of the foundations are essential as far the proper development, implementation, and use of these techniques.
@@
    Fully Observed There are no hidden variables, all quantities of interest are fully observed at every time step. Of course, graphs with hidden (latent) variables and entire structures are instrumental for modeling in many fields. However, the inclusion of latent variables and the required Expectation Maximization modification of the procedures described presents technical complications that add significant additional complexity, and thus would necessity a much greater length while obfuscating the message of the fundamentals.
We make a few departures from these assumptions throughout the work, which we explicitly indicate.
@@ 1.2 Applications of Dynamic Bayesian Networks
The discovery of dynamic Bayesian networks has found many applications, including medicine [Collett23, Eldawlatly08, lady22, Bueno16], economics [Ling15, LIU201946] and aviation [Matthews13, Valdes18, gomez18]. The applications related to aviation typically involve finding casual structure in a sub-problem on the dispatch of flights and focused on risk mitigation. The medical applications typically focus on either the discovery of fundamental principles concerning chemical reactions taking place in biological organisms or extracted clinical data. In economics, it is typically of interest to uncover relationships between the stock market and macroeconomic and other market indices that either influence or are influenced by it. In the following, we give some details about chosen applications as well as specific outcomes that modelling using DBNs has had in practice.
@@ 1.2.1 Medical Science
To highlight the importance of DBN-based discovery in medicine, we detail three separate applications. The first of these focuses on the quantification of disease development [Bueno16]. Understanding the progression of diseases is crucial in clinical medicine, as it informs the effectiveness of treatments. Most clinical medicine and pathology textbooks provide detailed descriptions of disease progression and treatment responses. However, there has been limited research quantifying these descriptions in detail. Typically, research examines the temporal aspect by describing treatment outcomes after a certain period. A significant challenge in gaining deeper insights is the relatively small size of clinical datasets, often comprising only a few hundred patients.
In the aforementioned contribution, a heuristic procedure is proposed for exploring and learning non-homogeneous time dynamic Bayesian networks, aiming to balance specificity and simplicity. The approach begins with a fully homogeneous (in time) model parts of which are gradually replaced with sub-models which reflect the expected structure change at a given level of time delay. Furthermore, a splitting technique is applied to further improve the predictive behavior of the model, such models are typically termed partitioned DBNs.
@@
The second application is concerned with the mapping of neural pathways [Eldawlatly08]. Identifying functional connectivity from simultaneously recorded spike trains is crucial for understanding how the brain processes information and instructs the body to perform complex tasks. The study investigates the applicability of dynamic Bayesian networks (DBNs) to infer the structure of neural circuits from observed spike trains. A probabilistic point process model was employed to assess performance. The results confirm the utility of DBNs in inferring functional connectivity as well as the directions of signal flow in cortical network models. Additionally, the findings demonstrate that DBNs outperform Granger causality when applied to systems with highly non-linear synaptic integration mechanisms.
The third chosen application focuses on the choice of appropriate treatment regimens for Chronic lymphocytic leukemia (CLL) [lady22]. This cancer is the most common blood cancer in adults, with a varied course and response to treatment among patients. This variability complicates the selection of the most appropriate treatment regimen and the prediction of disease progression. The aforementioned paper aims to develop and validate dynamic Bayesian networks (DBNs) to predict changes in the health status of patients with CLL and predict the progression of the disease over time. Two DBNs, the Health Status Network (HSN) and the Treatment Effect Network (TEN), were developed and implemented. Relationships linking the most important factors influencing health status and treatment effects in CLL patients were identified based on literature data and expert knowledge. The developed networks, particularly TEN, were able to predict the probability of survival in CLL patients, aligning with survival data collected in large medical registries. The networks can tailor predictions by integrating prior knowledge specific to an individual CLL patient. The proposed approach is a suitable foundation for developing artificial intelligence systems that assist in selecting treatments, thereby positively influencing the chances of survival for CLL patients.
@@ 1.2.2 Economics
The relationship between the stock market and national economies deepens as the market matures, highlighting the need to study their dynamic interplay. Economic indicators such as real income and savings rates play crucial roles in influencing stock market capitalization. Macroeconomic fundamentals wield considerable influence over both short and long-term periods. Some researchers argue that finance and economic growth are causally linked, suggesting the stock market’s potential to drive economic development. However, not all macroeconomic factors significantly impact stock prices. Understanding the strength of association among these variables offers insights into how the stock market behaves across varying economic landscapes. Research on the Chinese stock market examines how macroeconomic variables shape stock market indices over time, emphasizing the enduring influence of economic fundamentals amid short-term market fluctuations. The aforementioned interplay may be modeled by DBNs and has been detailed in [LIU201946].
@@ 1.2.3 Aviation
The global collection of aircraft and the airspace in which they operate is a complex system generating a vast amount of data, making it a challenging domain to model mathematically. This system includes critical elements such as aircraft, airports, flight crews, weather events, and routes, each with many subcomponents. For example, aircraft have numerous subsystems and components, each subjected to various stresses and maintenance actions, which influence their time dynamics. Airports have multiple runways and internal logistical processes, which influence the operational capacity. These system components interact in complex ways. For instance, each flight corresponds to an aircraft operated by a flight crew traveling from one airport to another via a route that may need to change due to weather. Multiple flights operate simultaneously, requiring coordination to avoid incidents while maximizing throughput and minimizing delays.
To give an idea about a specific aviation problem that may be tackled with DBNs, we describe the airport operation uncertainty characterisation developed in [gomez18]. The model outlines aircraft flow through the airport, emphasizing integrated airspace and airside operations. It characterizes various operational milestones based on an aircraft flow’s Business Process Model and Airport Collaborative Decision-Making methodology. Probability distributions for factors influencing aircraft processes need to be estimated, along with their conditional probability relationships. This approach results in a dynamic Bayesian network that manages uncertainties in aircraft operating times at the airport. The nodes of the network describe various aspects of the airport and flight operations. They cover meteorological conditions, arrival airspace variables such as timestamps and congestion metrics, airport infrastructure, operator and flight data, airside operational times and flight regulations, and the causes of delays.
@@ 2 Background - Dynamic Bayesian Networks
We present the general, and then specific forms, of DBN models. Consider that there is an $n_{x}$-dimensional stochastic process $X(t)$. The individual random variables $X_{i}(t)$ for all $i\in[n_{x}]$ can be valued as discrete, or as members of some field, such as $\mathbb{R}$. In addition, there can be an $n_{z}$-dimensional (time independent) random variable $Z$. Let us denote the generic spaces as $\mathcal{X}$ and $\mathcal{Z}$, respectively.
The defining character of DBNs is modeling the dependence of $X_{i}(t)$, for each $i$, on other quantities, i.e., defining the the evolution of the stochastic process $X(t)\to X(t+1)$. Formally, for the probability kernel defining the iterations of the stochastic process, the dependence must be Markovian, with additional time lagged auto-regressive effects, that is
@@
In addition, there is the important requirement that no in-time string of dependencies, that is from $X_{j\neq i}(t+1)$ to $X_{i}(t+1)$, forms a cycle. This presents the necessity of introducing graph theoretical notions to precisely characterize DBNs. Generically, we say a directed graph is a set of vertices $\mathcal{V}=\{v_{1},v_{2},...,v_{n}\}$ and edges =$\mathcal{E}=\{e_{1},e_{2},...,e_{m}\}$, where $e_{j}=(v_{j_{1}},v_{j_{2}})$ denotes the existence of a directed path between the two notes $v_{j_{1}}\to v_{j_{2}}$. We also say in this case that $j_{1}$ is a parent of $j_{2}$, or $j_{1}\in dpa(j_{2})$.
The DBN model structure is defined by a graph superset,
$\bar{\mathcal{G}}=\mathcal{G}(\mathcal{V}(X(t-1),X(t)),\mathcal{E}_{d})\cup\mathcal{G}(\mathcal{V}(X(t),X(t)),\mathcal{E}_{s})\cup\mathcal{G}(\mathcal{V}(X(t),Z),\mathcal{E}_{z})\cup\mathcal{G}(\mathcal{V}(X(t),X(t-\tau)),\mathcal{E}_{\tau})$ (2)
The first two union operands define connections in the model between the temporal random variables. That is $e=\{V_{1},V_{2}\}\in\mathcal{E}_{d}$ with $V_{i}\in\{\{X_{i}\}\}$ if the transition for $V_{2}=X_{i}(t+1)$, that is $p(X_{i}(t+1)\in A)$, is a function of $V_{1}=X_{j}$, or in other words $j\in dpa_{d}(i)$. For the static dependencies $j\in dpa_{s}(i)$, and with $V_{1}=X_{i}$ and $V_{2}=X_{j}$, it holds that $e=\{V_{1},V_{2}\}\in\mathcal{E}_{s}$ if $p(X_{j}(t+1)\in A)$ is a function of $V_{1}=X_{i}(t+1)$. Finally, we also have a (non-symmetric) matrix encoding the dependencies on the self-history $\tau\in dpa_{\tau}(i)\subset\{1,...,p\}$ and the dependencies on the static variables $dpa_{z}(i)=\left\{Z_{j}:\frac{\partial f_{i}(\cdot)}{\partial Z_{j}}\neq 0\right\}$. These, of course, can be encoded as graphs as well. Generically, the autoregressive effects can include lagged influence across distinct random variables, however for simplicity we focus only on including self-dependencies for time lags longer than one.
@@
See Figure 1 for an illustration of a DBN and its transition dynamics. In this case for $i=1$, the Markovian transitions are from itself and from $X_{2}$, and there are no intra time nodes or static nodes directed to it, and so $V_{dpa(1)}(t)=V_{dpa_{d}(1)}=\{X_{1}(t-1),X_{2}(t-1)\}$. For node 2, there are no Markovian transitions and two intra-node dependencies, thus $V_{dpa(2)}(t)=V_{dpa_{s}(2)}(t)=\{X_{1}(t),X_{3}(t)\}$. Finally, for $X_{3}(t)$, there are two Markovian dependencies, and a static covariate dependence. Thus $V_{dpa(3)}(t)=V_{dpa_{d}(3)\cup dpa_{z}(3)}(t)=\{X_{1}(t-1),X_{3}(t-1),Z\}$.